Four wooden rods with length of 70cm, 119cm, 84cm and 105cm are cut into pieces of the same length.find the greatest possible length for these pieces if no wood is left over.
Answer by josgarithmetic(39617)

Each of the four numbers is divisible by 7.
No other factor common to all of them.



-
70=10*7,119=7*17,84=12*7,105=3*5*7